OpenStudy (anonymous):

What is the solution for: 3x+2y=6 X+3y=9

OpenStudy (anonymous):

the first one we got: 2y=6-3x y= (6-3x)/2 y=3-(3/2)x the second one we got: 3y=9-x y=(9-x)/3 y=3-x/3 Now we get (1)=(2) so: 3-(3/2)x=3-x/3 solve this and u'll get the value of x, then find the value of y